Consider the market for Relaxo floaters where the demand curve is given by Q ^ d = 200 - 2P and Q^ prime =250+30P^ prime -10 P^ m . where Q ^ d is the quantity demanded, Q ^ 1 is quantity supplied. P ist the price of Relaxo floaters.




i. What is the maximum price at which the consumers are willing to purchase the floaters?




Calculate the price elasticity of demand and supply at equilibrium.




11. If now a unit tax is imposed on floaters, which side (demand or supply) will bear a higher




burden of the tax? Why?




Show that the demand side is more elastic than supply as long the price is above 50.

1
Expert's answer
2022-03-25T15:07:36-0400

i) What is the maximum price at which the consumers are willing to purchase the floaters?

Here, we equate the demand and supply equations

"200 - 2P =" "250+30P -10 P^2"

"250-200=-2P-30P+10P^2"

"50=-32P+10P^2"

"10P^2-32P-50=0"

"P=4.35"


Calculate the price elasticity of demand and supply at equilibrium.

"Q=200-2(4.35)"

"Q=191.3"

Elasticity of demand at equilibrium is the same as the elasticity of supply.

"e=\frac{\Delta Q }{\Delta P} =43.97"


 If now a unit tax is imposed on floaters, which side (demand or supply) will bear a higher

The demand side will bear the tax burden since the payment of tax will be passed from the supplier through the retailers to the consumer and they will pay the taxes through VAT


Show that the demand side is more elastic than supply as long the price is above 50

"Q = 200-2(50) = 100"

"e = \frac{100}{50}= +2"